dy/dx=x+3y-5/x-y-1


1
Expert's answer
2021-02-24T06:32:02-0500

"\displaystyle\frac{dy}{dx}=\frac{x+3y-5}{x-y-1}"

Solution:

Introduce replacements: "x=x_0+t" and "y=y_0+z" , where "x_0" and "y_0" are solutions of the system of equations:

"\begin{cases}
 x_0+3y_0-5=0 \\
 x_0-y_0-1=0 
\end{cases}"

"x_0=2" , "y_0=1" .

Substitute "x=2+t" and "y=1+z" into the initial equation:

"\displaystyle\frac{dy}{dx}=\frac{\frac{dy}{dt}}{\frac{dx}{dt}}=\frac{dz}{dt}=\frac{2+t+3(1+z)-5}{2+t-(1+z)-1}"

"\displaystyle\frac{dz}{dt}=\frac{t+3z}{t-z}"

Now introduce another replacement: "z=ut" . Then "\frac{dz}{dt}=\frac{du}{dt}t+u" .

"\displaystyle\frac{du}{dt}t+u=\frac{t+3ut}{t-ut}"

"\displaystyle\frac{du}{dt}t=\frac{1+3u}{1-u}-u=\frac{1+3u-u+u^2}{1-u}=\frac{(1+u)^2}{1-u}"

"\displaystyle\frac{1-u}{(1+u)^2}du=\frac{dt}{t}"

"\displaystyle-\frac{(u+1-2)}{(1+u)^2}du=\frac{dt}{t}"

"\displaystyle(-\frac{1}{1+u}+\frac{2}{(1+u)^2})du=\frac{dt}{t}"

"\displaystyle-\ln|1+u|-\frac{2}{1+u}=\ln{|t|}+\ln{C}"

"\displaystyle(1+u)\cdot t\cdot C=e^{-\frac{2}{1+u}}" .

Returning to replacements: "z=y-1" , "t=x-2" and "\displaystyle u=\frac{z}{t}=\frac{y-1}{x-2}"

"\displaystyle(1+\frac{y-1}{x-2})(x-2)\cdot C=e^{-\frac{2}{1+\frac{y-1}{x-2}}}"

"\displaystyle(x+y-3)\cdot C=e^{-\frac{2(x-2)}{x+y-3}}"

Answer: "\displaystyle(x+y-3)\cdot C=e^{-\frac{2(x-2)}{x+y-3}}" .